Two boats are heading away from shore. Boat 1 heads due north at a speed of 4 m/s relative to the shore. Relative to boat 1, boat 2 is moving 30° north of east at a speed of 1.2 m/s. A passenger on boat 2 walks due east across the deck at a speed of 1.1 m/s relative to boat 2. What is the speed of the passenger relative to the shore?
Answer #1

here,

the speed of boat 1 , v1 = 4 m/s j

the speed of boat 2 relative to boat 1 , v21 = 1.2 m/s * ( cos(30) i + sin(30) j)

v21 = 1.04 i m/s + 0.6 j m/s

the speed of passenger relative to boat 2 , v2 = 1.1 m/s i

the velocity of the passenger relative to the shore , v = v1 + v2 + v3

v = 4 j + 1.04 i + 0.6 j + 1.1 i

v = 2.14 i m/s + 4.6 j m/s

speed of the passenger relative to the shore, |v| = sqrt(2.14^2+ 4.6^2) = 5.07 m/s